Find the amount that results from the given investment.
$20 invested at 8% compounded continuously after a period of 2 years
After 2 years, the investment results in $?
(round to the nearest cent as needed)

Answer by mathmate(429)

Question:
Find the amount that results from the given investment.
$20 invested at 8% compounded continuously after a period of 2 years
After 2 years, the investment results in $?
(round to the nearest cent as needed)

Solution:
P=$20=principal
i=8% annual continuous
t=2 years

The accumulated amount is
F=Pe^(it)

Therefore after two years, accumulated amount
F=20*e^(0.08*2)
=20*1.1735
=$23.47